OpenStudy (dls):

If a,c,b are in GP,then the line ax+by+c=0 a)has a fixed direction b)always passes through a fixed point c)forms a triangle with the axes whose area is constant d)always cuts intercepts on the axes such that their sum=0

OpenStudy (dls):

geometric progression

OpenStudy (hba):

good question :P @skullpatrol

OpenStudy (dls):

lol

OpenStudy (shubhamsrg):

b=ar c=ar^2 => x + ry + r^2 =0 x/(-r^2) + y/(-r) = 1 so ? constant area ?

OpenStudy (dls):

right

OpenStudy (dls):

but what did u do O.o

OpenStudy (shubhamsrg):

do it like this : x intercept is -c/a y intercept is -c/b area = 1/2 (c^2 /ab) and they are in GP,,and hence your answer..

OpenStudy (shubhamsrg):

a,c,b hai,,a,b,c, nahi..sorry,,a,c,b lene se intercept wale method se bhi aa jayega.. maine waha constant isliye bol diya tha kyunki socha kisi r ke liye constant hoga! :P